# Centimeter-sized Objects at Micrometer Resolution: Extending Field-of-View in Wavefront Marker X-ray Phase-Contrast Tomography

### v1

@@ 3.3 Measurement Protocol
The rat brain is imaged using a hanging-rotation-axis approach, as illustrated in Figure 3. The sample is placed inside a tube with a diameter of and attached to a rotation stage at the top. To minimize variations in electron density, the tube is submerged in a bath filled with ethanol, which prevents the strong changes in phase that would occur if measured against air. This setup also maintains a consistent dynamic range for the phase retrieval method across all scans, which would not be the case if some regions contain air while the inner region is fully occupied by the sample. At the high radiation doses encountered at synchrotron-radiation sources, ethanol is preferred over water because water is prone to radiolysis. This leads to the formation of gas bubbles that can displace tissue during measurements and increase phase variations.
@@
The scans are recorded at an energy of using a continuous rotation mode; the position of the wavefront marker is moved after each 180∘ rotation, resulting in 9 steps per scanned FOV. At the beginning and the end of each scan a set of 70 flat-field images are taken. In total, 4000 projections are taken with an exposure time of each.
3.4 Algorithm Parameters
@@
The images are downsampled with a factor of 8 to increase the computational efficiency and mitigate the influence of the speckle background on the registration result. This mitigation is performed since the speckle background inside the overlapping region is not necessarily the same for the two scans and may therefore hinder the registration algorithm from matching the actual sample features. Considering a tradeoff between computational efficiency and deformation complicity, 16 control points on a grid are assigned at the overlapping region. To emphasize on the local motion, the local normalized cross-correlation [31] is used as the similarity metric for the registration.
@@
After the registration, the optimized deformation is applied to the image with the original resolution. The deformed image is then passed to the further steps. The geometrical error induced by beam divergence is neglected given the approximately parallel beam encountered at most synchrotron radiation sources.
@@
For computational efficiency, only a subregion of the overlapping regions pixels in size is used for the optimization. Since the target function wraps the phase images that are dynamically retrieved via the phase retrieval optimization, the target function is non-differentiable. We therefore use the derivative-free optimizer BOBYQA [32] from the NLopt library [33]. The optimization variables are the weights of eigenflats, they are therefore unitless. The boundary is empirically set to , the initial step is 0.1 and the optimization is terminated when the relative change of the cost value falls below . We use the first three components (eigenflats) for the flat-field generation, which results in a total of 27 optimization variables ( weights steps).
4 Results and Discussion
@@
Figure 4 shows a subregion of a projection taken during the inner scan. More specifically, the differential phase-contrast signal in the -direction is shown. It is once obtained by a conventional flat-field correction using a mean of the 70 recorded flat-field (Figure 4a) images and by our optimization approach (Figure 4b). While in Figure 4a, remnants of the 2D grating structure are superimposed onto the anatomical information, the effect is strongly reduced in Figure 4b by the correction. The same correction is repeated for all 4000 projection images of the middle part of the scan. The displacements of the FFD control points as determined by the registration are below 10 pixels. This marks the maximum amount by which image pixels at the overlapping region are shifted. The three subregions of the sample are then stitched as previously described to obtain a larger combined FOV (see Figure 4c).
@@
After the 2D phase retrieval using UMPA with a window size of 3 and reconstruction of the 2D Fourier integrated phase information using a filtered back projection, an electron density map displaying the entire horizontal extent of the rat brain is obtained. Figure 5 shows an axial slice of the brain. As becomes apparent, no obvious ring artifacts are visible in the scan. These would be created by grating structures in the projections. Moreover, no obvious artifacts due to the stitching of the three scans are present. At the same time, slight changes in electron density throughout the brain become discernible. The measured electron density is similar to the value of 305 previously reported for a human cerebellum [34].
@@
The overlap in this scan encompassed 1183 pixels in the horizontal direction, but may be reduced by approximately a factor of 6 in size in future scans. This is because only a subregion 180 pixels in horizontal extent of the overlap is used for reference-image matching. Such a reduction in overlap width is desirable to prevent redundancy in the acquired data leading to unnecessarily prolonged measurement times.
The spatial resolution for the rat brain scan is calculated on a slice of the tomogram using a Fourier Shell Correlation (FRC) method [35, 36] with four image subsets. The resolution is determined to be according to the full-bit and according to the half-bit criterion. This resolution is most likely not the optimum achievable value since the sample was undersampled significantly in this proof-of-principle experiment. Given the approximately 5000 pixels (for 2-fold binning) inside the tube walls, 8000 projection angles would be required according to the Nyquist sampling theorem, but only 4000 projection angles were acquired. A higher number of angles should therefore be chosen for future experiments to obtain the highest possible resolution.
@@
In the results of the proposed eigenflat optimization shown in Figure 4b, a residual is partially visible as diagonal patterns. This may be attributed to the fact that due to efficiency considerations we do not use all potentially available eigenflats, meaning not all the potentially available eigenflat-field information is considered for the optimization.
5 Conclusion and Outlook
@@
In conclusion, we have presented a novel approach significantly expanding the application range of wavefront-marker-based imaging techniques. Our method enables the retrieval of small differences in electron density for centimetre-sized objects, given sufficient transmission, by cascading beam reference information inward from outer scans and combining subscans using deformable registration. This advancement is particularly relevant for upcoming fourth-generation synchrotron sources, which will feature smaller fields of view, and enables hierarchical scans yielding electron density maps comparable across beamlines, facilitating collaboration and validation of results. The ability to obtain quantitative electron density information for larger samples is a significant step forward in phase-contrast imaging, as it opens up new possibilities for studying complex biological structures and materials at multiple scales. This could be particularly interesting for fields such as biomedical imaging, materials science, and paleontology, where understanding the details of large objects is crucial.
@@
The method was demonstrated using speckle-based imaging but is in principle applicable to any phase-contrast or attenuation-contrast method requiring accurate flat-field information, provided a suitable optimization metric is available. As described in the supplementary document, the method also allows for measurements of even larger objects with Talbot Array Illuminators using the cascading approach. However, it remains to be experimentally investigated whether a large number of stitchings would lead to an accumulation of error when propagating information from the outermost scans inward. In such a case however, one may consider correcting the reference information locally for each FOV instead of using the overlap regions, as demonstrated for the ROI scan in the supplement.
@@
Another drawback is the computational intensity of the optimization procedure for each pair of overlapping projections. Future work could explore adapting machine learning-based approaches for synthetic reference generation, such as those developed by [37], to alleviate this issue for the case of wavefront-marker based phase-contrast imaging as well.
[…]
Acknowledgments The authors thank Fabio De Marco for fruitful discussions and Natalia S. Pellegata from Helmholtz Munich for providing the rat brain sample.
@@
Disclosures Junan Chen and Matthias Wieczorek are employees of ImFusion GmbH. The remaining authors declare no conflicts of interest.
@@
Data availability Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.

### v2

@@ III-C Measurement protocol
The rat brain is imaged using a hanging-rotation-axis approach, as illustrated in Figure 3. The sample is placed inside a tube with a diameter of and attached to a rotation stage at the top. To minimize variations in electron density, the tube is submerged in a bath filled with ethanol, which prevents the strong changes in phase that would occur if measured against air. This setup also maintains a consistent dynamic range for the phase retrieval method across all scans, which would not be the case if some regions contain air while the inner region is fully occupied by the sample. At the high radiation doses encountered at synchrotron-radiation sources, ethanol is preferred over water because water is prone to radiolysis. This leads to the formation of gas bubbles that can displace tissue during measurements, thereby increasing phase variations.
@@
The scans are recorded at an energy of using a continuous rotation mode; the position of the modulator is moved after each 180∘ rotation for a total of 9 stepping positions per scanned FOV. At the beginning and end of each scan, a set of 70 flat-field images is taken. In total, 4000 projections are acquired for each scan with a -exposure time. In this work, a complete tomogram is first acquired for the left FOV, as depicted in Figure 1, followed by a complete tomogram for the inner and right FOV. Including motor overhead for the modulator movement, the total scan time is approximately .
III-D Algorithm parameters
@@
The images are downsampled with a factor of 8 to increase the computational efficiency and mitigate the influence of the modulator background on the registration result. The factor was empirically determined to be sufficient for the modulator pattern to disappear. This is required since the modulation background inside the overlapping region is not necessarily the same for the two scans and may therefore hinder the registration algorithm from matching the actual sample features. Considering a tradeoff between computational efficiency and deformation complexity, 16 control points on a grid are assigned at the overlapping region, resulting in 32 degrees of freedom. To emphasize the local motion, the local normalized cross-correlation [3] is used as the similarity metric for the registration.
@@
Note that 8-fold downsampling is applied only during the registration step. After registration, the optimized deformation is applied to the image at the original resolution before the 8-fold downsampling. The deformed image is then passed to the further steps. The geometrical error induced by beam divergence is neglected, given the approximately parallel beam encountered at most synchrotron radiation sources.
@@
For computational efficiency, only a subregion of the overlapping regions pixels in size is used for the optimization. In our experiments, larger regions did not lead to significant improvements in reference matching, while introducing additional computational complexity. Since the target function wraps the phase images that are dynamically retrieved via the phase retrieval optimization, the target function is non-differentiable. We therefore use the derivative-free optimizer BOBYQA [26] from the NLopt library [16]. The optimization variables are the weights of eigenflats, they are therefore unitless. The boundary is empirically set to , the initial step size is 0.1 and the optimization is terminated when the relative change of the cost value falls below . We use the first three components (eigenflats) for the flat-field generation, which results in a total of 27 optimization variables ( weights steps).
IV Results and discussion
We ran the image processing pipeline on the high-performance computing cluster of DESY (Hamburg, Germany), with the approximate runtimes for each processing step in Figure 2 stated in the following: The matching reference images were calculated separately for each projection in the left and right FOV (Figure 2(a)), taking per projection. The phase retrieval took a further per projection. Both operations were performed concurrently on 100 compute nodes, each equipped with an AMD EPYC 75F3 CPU. The deformable registration (Figure 2(b)) was performed on a single GPU node equipped with an NVIDIA Quadro RTX 8000 and took per projection pair. The calculation of matching reference images using the adapted eigenflat approach for the middle projections (Figure 2(c)) was again performed in parallel on 100 CPU nodes and required per projection, followed by UMPA phase retrieval with the same runtime as stated for the left and right FOV.
Figure 4 shows a subregion of a projection taken during the inner scan, specifically the differential phase-contrast signal in the -direction. The image is obtained once by a conventional flat-field correction using the mean of the 70 recorded flat-field images (Figure 4(a)) and once by our optimization approach (Figure 4(b)). While remnants of the 2D grating structure are superimposed onto the anatomical information in (a), this effect is strongly reduced in (b) by the optimized correction. To provide a quantitative comparison of the image quality, we compute the standard deviation and TV within the ROI indicated by the white dashed box. The standard deviation is reduced from 0.067 to 0.062 (), while the total variation is reduced from 423.01 to 330.57 (). These improvements in both metrics align with the visual evaluation. Note that reducing these metrics to zero is not desired, as this would indicate the removal of actual sample features.
@@ IV Results and discussion
The same correction is repeated for all 4000 projection images of the middle part of the scan. The displacements of the FFD control points as determined by the registration are below . This represents the maximum displacement of image pixels in the overlapping region. The three subregions of the sample are then stitched as previously described to obtain a larger combined FOV, as depicted in Figure 4(c).
@@
After 2D phase retrieval using UMPA with a window size of 3 and reconstruction of the 2D Fourier-integrated phase information using filtered backprojection, an electron density map displaying the entire horizontal extent of the rat brain is obtained. Figure 5 shows an axial slice of the brain. As becomes apparent, no obvious ring artifacts are visible in the scan, which would otherwise be created by residual grating structures in the projections. Moreover, no obvious artifacts due to the stitching of the three scans are present. At the same time, slight changes in electron density throughout the brain are clearly discernible. The measured electron density is similar to the value of previously reported for a human cerebellum [38].
@@
The overlap in this scan encompassed 1183 pixels in the horizontal direction, but may be reduced by approximately a factor of 6 in size in future scans. This is because only a subregion 180 pixels in horizontal extent of the overlap is used for reference-image matching. Such a reduction in overlap width is desirable to prevent redundancy in the acquired data, leading to unnecessarily prolonged measurement times.
The spatial resolution for the rat brain scan is calculated on a slice of the tomogram using a Fourier ring correlation (FRC) method [39, 43] with four image subsets [30]. The resolution is determined to be according to the full-bit and according to the half-bit criterion. A resolution analysis on different regions of the image confirmed a spatially homogeneous resolution. The final resolution of the scan is affected by multiple parameters: One is the UMPA window parameter, where large windows lead to better noise statistics at the cost of a lower resolution. While the experiment could have been conducted without 2-fold binning of the data, this would have led to a significant increase in computational complexity and scan time due to the additional angles required by the Nyquist sampling theorem.
@@
In the results of the proposed eigenflat optimization shown in Figure 4(b), a residual is partially visible as diagonal patterns. This may be attributed to the fact that, due to computational efficiency considerations, we do not utilize all potentially available eigenflats, meaning that not all potentially available eigenflat-field information is considered for the optimization.
V Conclusion and Outlook
@@
In conclusion, we have presented a novel approach that significantly expands the application range of modulation-based imaging techniques. Our method enables the retrieval of small differences in electron density for centimeter-sized objects, given sufficient transmission, by cascading beam reference information inward from outer scans and combining subscans using deformable registration. This advancement is particularly relevant for fourth-generation synchrotron sources, which will feature smaller fields of view, and enables hierarchical scans yielding electron density maps comparable across beamlines, facilitating collaboration and validation of results. The ability to obtain quantitative electron density information for larger samples is a significant step forward in phase-contrast imaging, as it opens up new possibilities for studying complex biological structures and materials at multiple scales. This could be particularly interesting for fields such as biomedical imaging, materials science, and paleontology, where understanding the details of large objects is crucial.
The method was demonstrated for modulation-based imaging but is applicable in principle to any phase-contrast or attenuation-contrast method requiring accurate flat-field information, provided a suitable optimization metric is available. In this work, we successfully used the total variation (TV) as the optimization metric. In the future, more metrics beyond the native TV should be investigated to simultaneously capture the potential residual low-frequency mismatches in the flat-field images that may occur due to larger-scale shifts of the illumination. While TV is well-suited for characterizing residual modulator patterns in overlap regions, adaptations may be required for region-of-interest scans when samples exhibit pronounced high-frequency spatial components, as the TV metric may inadvertently penalize genuine sample features.
@@
As described in the appendix, our method also allows for measurements of even larger objects than shown here using the cascading approach. However, experimental investigation is needed to determine whether a large number of stitchings would lead to an accumulation of error when propagating information from the outermost scans inward. In such a case, however, one may consider correcting the reference information locally for each FOV instead of using the overlap regions, as demonstrated for the ROI scan in the appendix. For sub-scans, there will be overlap regions that require reference matching, leading to increased computational cost. The precise amount of overlap between scans required likely depends on the experimental conditions and the modulator design, and will therefore need to be determined empirically.
@@
A drawback of our method is the computational intensity of the optimization procedure for each pair of overlapping projections. Future work could explore adapting machine learning-based approaches for synthetic reference generation, such as those developed by Ref. [12], to alleviate this issue for modulation-based phase-contrast imaging as well.
In general, projection stitching approaches require relatively stable samples [5, 44], as structural changes or movements larger than the pixel size during acquisition can cause irrecoverable artifacts [6]. This limitation is particularly problematic for fresh biological tissues that undergo degradation during extended scan times. If sub-volumes can be captured quickly enough compared to the timescale of the sample changes, a volume stitching approach offers the advantage of capturing multiple self-consistent snapshots of the sample [6, 45].
Additionally, since a main advantage of modulation-based imaging is that it does not rely on the single-material assumption, future studies should examine samples that strongly violate this assumption to further demonstrate the method’s capabilities. Examples include imaging metals embedded in organic tissue [30] or other heterogeneous materials. Furthermore, since modulation-based imaging provides a dark-field signal [49, 37], another promising avenue for future research is examining this contrast mechanism for larger samples.
Acknowledgments
The authors gratefully acknowledge financial support by the ERC Consolidator Grant (Julia Herzen, TUM, DEPICT, PE3, 101125761) and by the EIC Pathfinder (1MICRON, 101186826). D.J. acknowledges a doctoral scholarship by the Friedrich Naumann Foundation for Freedom funded by the Bundesministerium für Forschung, Technologie und Raumfahrt (BMFTR). We acknowledge DESY (Hamburg, Germany), a member of the Helmholtz Association HGF, for the provision of experimental facilities. Parts of this research were carried out at PETRA III and we would like to thank Fabian Wilde for assistance in using the Beamline P05, operated by the Helmholtz-Zentrum Hereon. This research was supported in part through the Maxwell computational resources operated at DESY. The authors thank Fabio De Marco for fruitful discussions and Natalia S. Pellegata from Helmholtz Munich for providing the rat brain sample.
Disclosures
@@ Acknowledgments
Junan Chen and Matthias Wieczorek are employees of ImFusion GmbH. The remaining authors declare no conflicts of interest.
Data and code availability
@@
Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request. A Python implementation of the algorithm is also available upon request.
Appendix 1: Cascading reference approach
We perform a proof-of-concept experiment to verify the feasibility of the cascading approach mentioned in Section II-F. In this approach, information from the outermost projection is propagated inward to the innermost projection. To test the procedure, we use a measurement that contains three overlapping scans, as shown in Figure 1(b), but do not correct the left outer projection using the conventional approach (i.e., a dynamic flat-field correction using a sample-free reference region). Instead, we assume the left projection to be a scan without a sample-free reference area and use information from the left overlap region, depicted in Figure 1(b), to find the optimal flatfield. After correcting the inner scan as shown in Figure 4(b), we find the optimal flatfield for the leftmost scan using the optimization term from Eq. (9). Figure 6 shows a comparison between the resulting left outer phase image corrected using (a) the mean flatfield and (b) the optimized flatfield obtained via our cascading approach. As becomes apparent, the artifacts caused by leftover modulator structures are reduced in (b). This demonstrates that our approach is also feasible for 3 inner regions, as would be encountered for regions to be stitched.
Appendix 2: Sandpaper modulator
Since the method is applied to Talbot array illuminators in the main section of this work, we want to demonstrate that its working principle is also applicable to sandpaper, a different type of commonly used modulator in modulation-based imaging.
For this purpose, projection data previously acquired in Ref. [34] with an effective pixel size of using a very similar setup is used. The examined object is a phantom sample consisting of polymethylmethacrylate (PMMA), polytetrafluorethylene (PTFE), polychlorotrifluoroethylene (PCTFE), and polyamide nylon-(6,6). It also includes a toothpick and a pipette tip containing polystyrene microspheres.
The modulator used for this scan is P1000 sandpaper with a mean particle size of . The same approach as mentioned in Section Appendix 3: Region-of-interest scans is applied to the acquired data. Figure 7 shows a ROI of the differential phase image retrieved with a mean flatfield (a) and with an optimized flatfield from our method (b). In Figure 7(b), a reduction in background noise is observed, which marks the removal of the random speckle pattern generated by the sandpaper modulator. Moreover, the standard deviation of the differential phase image decreases from 0.051 to 0.034 (). The result visually and quantitatively confirms the feasibility of the proposed method on data acquired with a sandpaper modulator.
Appendix 3: Region-of-interest scans
The reference region used in the proposed optimization is not always available in certain cases, an example being ROI scans. However, a similar strategy is still feasible for the purpose of dynamic flat-field correction. One can adapt the proposed method by evaluating a sample region only. The optimization problem is then turned into minimizing the cost function
where denotes a small region of interest inside the sample from the differential phase image. A ROI with a relatively low variance is selected in order to mitigate the influence of the sample structures on the TV evaluation. Figure 8 provides a comparison of the result from the proposed eigenflat-optimization and the other approaches. Note that the scan itself is not originally a ROI scan since the edge of the sample has been imaged as well, but the behavior of the eigenflat method is simulated by considering a ROI inside the sample region. At the same time, a comparison with approaches using the sample-free reference area becomes possible.
Figure 8(a) is the phase image acquired with a mean flatfield. Panel (b) shows the result using the method proposed in Ref. [29]. Note that this method is not applicable for ROI scans given that a sample-free reference region is required, so the result is only provided for comparison. Panel (c) shows the result from the same method as in Figure 8(b), using a sample region instead. Figure 8(d) is the phase image retrieved using the flatfield generated from the proposed optimization in Eq. (10). In analogy to Figure 8(c), a sample region is used for the optimization. Our result in Figure 8(d) shows a comparable quality to the result in Figure 8(b), however, our method does not require a sample-free reference region as in the other method. Moreover, fewer eigenflat components are needed for our method to generate an image with similar quality (3 vs. 5).